OpenStudy (anonymous):

For the equations 1:x(L-1)=0, 2: y(L+1)=0, 3: x^2+y^2=4. Why does the solution say "L=1 gives y=0" and "L=-1 gives x=0", how do I find out that myself?

OpenStudy (anonymous):

Let 1=2 Set L=1 to get zero on the left hand side, you'll then have x*0=y*2 which gives y is always 0 (because of x*0) Set L=-1 to get zero on the right hand side, you'll then have x*-2=y*0 which gives x is always 0 (because of y*0)